OpenStudy (anonymous):

find the radius of a circle whose circumference is equal to the sum of the circumference of two circles with radii of 4 inches and 6 inches

OpenStudy (ajprincess):

Circumference of a circle=\(2\pi r\) \(2\pi r=2\pi*4+2\pi*6\) Can u find r nw @fangmitchell?

OpenStudy (anonymous):

62.8?

OpenStudy (anonymous):

answer=10

OpenStudy (ajprincess):

hmm no. \(2\pi r=2\pi*4+2\pi*6\) Factor out \(2\pi\) from the right habd side \(2\pi r=2\pi(4+6)\) \(2\pi r=2\pi(10)\) Divide \(2\pi\) both sides. \(r=10inches\) Is it clear @fangmitchell?

OpenStudy (ajprincess):

*hand

OpenStudy (anonymous):

oh yes i get it! thanks :)

OpenStudy (ajprincess):

yw:)